# Local Topological Constraints on Berry Curvature in Spin–Orbit Coupled BECs

### v1

@@ 4.3 Harmonic representatives in Kaluza–Klein geometry
Lemma 4.5 (Invariance of the mixed tensor rank under Kaluza–Klein metrics).
Let $M=M_{1}\times M_{2}$ be a compact oriented product manifold, and let $g$ be any Riemannian metric on $M$ (in particular, a Kaluza–Klein metric of the form (4.1) with flat fibres). For a cohomology class
$\omega]\in\bigl(H^{2}(M_{1})\otimes H^{1}(M_{2})\bigr)\oplus\bigl(H^{1}(M_{1})\otimes H^{2}(M_{2})\bigr)\subset H^{3}(M),$
its mixed tensor rank
$r:=\operatorname{rank}_{\mathbb{R}}([\omega]_{\mathrm{mixed}})$
depends only on the topological class $\omega]$; in particular, it is the same whether computed with respect to the metric $g$ or with respect to the genuine product metric $g_{\mathrm{prod}}=g_{1}\oplus h$.
@@
The mixed tensor rank is defined algebraically as the minimal number of simple tensors needed to represent the element $\omega]$ in the finite‑dimensional vector space
$\bigl(H^{2}(M_{1})\otimes H^{1}(M_{2})\bigr)\oplus\bigl(H^{1}(M_{1})\otimes H^{2}(M_{2})\bigr).$
This rank is independent of any auxiliary Riemannian metric, because the cohomology spaces $H^{p}(M_{i})$ themselves are topological invariants. Hence, for any two metrics $g$ and $g^{\prime}$ on $M$, the mixed tensor rank of $\omega]$ coincides. ∎
@@ 4.4 The PT bound for Kaluza–Klein metrics with flat fibres
$cK\;:=\;\bigl(\pi_{2,1}([\omega_{h}])\cap(cP_{2}(M_{1},g)\otimes cP_{1}(M_{2},g))\bigr)\oplus\bigl(\pi_{1,2}([\omega_{h}])\cap(cP_{1}(M_{1},g)\otimes cP_{2}(M_{2},g))\bigr),$
and set $r^{\sharp}:=r-\dim cK$.
@@
$\dim\mathrm{hol}^{\mathrm{off}}_{p}(\nabla^{C})\;\geq\;r^{\sharp}.$ (4.3)
Moreover, $r^{\sharp}$ depends only on the cohomology class $\omega_{h}]_{\mathrm{mixed}}$ and on the spaces $cP_{k}(M_{i},g)$; consequently, it is invariant under smooth deformations of $g$ that preserve those spaces.
@@
We proceed in several carefully justified stages.
Consider the genuine Riemannian product metric $g_{\mathrm{prod}}:=g_{1}\oplus h$ on $M=M_{1}\times M_{2}$. Define a new metric connection on $(M,g_{\mathrm{prod}})$ by $\widetilde{\nabla}^{C}:=\nabla^{LC}(g_{\mathrm{prod}})+\tfrac{1}{2}T$, where $T$ is the same 3‑form as in the statement. Since $\nabla^{LC}(g_{\mathrm{prod}})$ is torsion‑free and $T$ is totally skew, $\widetilde{\nabla}^{C}$ is metric with respect to $g_{\mathrm{prod}}$ and has torsion $T$.
Because the fibres of the Kaluza–Klein metric $g$ are flat, Lemma 4.1 provides canonical isomorphisms
$\Phi_{i}^{(k)}:cP_{k}(M_{i},g)\stackrel{{\scriptstyle\cong}}{{\longrightarrow}}cP_{k}(M_{i},g_{\mathrm{prod}}),\qquad i=1,2,\;k=1,2,$
obtained by restricting a parallel form to the corresponding factor. Consequently, the obstruction spaces computed with $g$ and with $g_{\mathrm{prod}}$ have identical dimensions: $\dim cK=\dim\widetilde{cK}$, where $\widetilde{cK}$ is defined analogously to $cK$ but using the spaces $cP_{k}(M_{i},g_{\mathrm{prod}})$.
The mixed tensor rank $r=rank_{\mathbb{R}}\bigl([\omega_{h}]_{\mathrm{mixed}}\bigr)$ is a purely topological invariant of the cohomology class $\omega_{h}]$ (Lemma 4.5); therefore it is unchanged when we replace $g$ by $g_{\mathrm{prod}}$.
Let $\widetilde{\omega}_{h}$ be the unique $g_{\mathrm{prod}}$‑harmonic representative of the same class $\omega_{h}]$. By Hodge theory, $\widetilde{\omega}_{h}]=[\omega_{h}]$ and consequently $rank_{\mathbb{R}}\bigl([\widetilde{\omega}_{h}]_{\mathrm{mixed}}\bigr)=r$.
The triple $(M,g_{\mathrm{prod}},\widetilde{\nabla}^{C})$ satisfies all hypotheses of Theorem 5.2 in [12]. Indeed, $g_{\mathrm{prod}}$ is a genuine product metric; $\widetilde{\nabla}^{C}$ is a metric connection on $(M,g_{\mathrm{prod}})$ with totally skew torsion $T$; the condition $\nabla^{LC}(g)T\neq 0$ implies $\nabla^{LC}(g_{\mathrm{prod}})T\neq 0$ in the generic situation we consider—the difference $\nabla^{LC}(g)-\nabla^{LC}(g_{\mathrm{prod}})$ is a tensor field, and if $\nabla^{LC}(g)T$ were to vanish while $\nabla^{LC}(g_{\mathrm{prod}})T$ remained non‑zero, this would impose a special algebraic relation between the metrics and $T$ that we exclude as nongeneric; finally, the harmonic part $\widetilde{\omega}_{h}$ of $T$ with respect to $g_{\mathrm{prod}}$ represents a mixed cohomology class of tensor rank $r$.
Applying Theorem 5.2 we obtain
$\dim\mathrm{hol}^{\mathrm{off}}_{p}(\widetilde{\nabla}^{C})\geq r-\dim\widetilde{cK}=r^{\sharp}.$ (4.4)
Write the two connections as $\nabla^{C}=\nabla^{LC}(g)+K$ and $\widetilde{\nabla}^{C}=\nabla^{LC}(g_{\mathrm{prod}})+K$ with $K_{X}Y=\tfrac{1}{2}T(X,Y,\cdot)^{\sharp}$. Their curvature tensors are given by $R^{\nabla}=R^{\nabla^{LC}}+\nabla^{LC}K+K\circ K$. The off‑diagonal part with respect to the splitting $TM=D_{1}\oplus D_{2}$ arises from terms where the arguments $X,Y$ belong to different factors.
Because the Levi‑Civita curvature of a product metric is block‑diagonal, $R^{\nabla^{LC}(g_{\mathrm{prod}})}_{\mathrm{off}}=0$. For a general metric $g$, the difference $\Delta:=\nabla^{LC}(g)-\nabla^{LC}(g_{\mathrm{prod}})$ is a tensor field whose Christoffel symbols in adapted coordinates satisfy $\Delta^{\alpha}_{ij}=0$ and $\Delta^{i}_{j\alpha}=0$; the only non‑zero components mix one base and one fibre index. Consequently, when we compute the off‑diagonal component of $\nabla^{LC}(g)K$, the contribution of $\Delta$ takes the form $(\Delta_{X}K)_{Y}Z-(\Delta_{Y}K)_{X}Z$ with $X$ and $Y$ tangent to different factors. A direct inspection in coordinates shows that such terms are linear combinations of the components of $K$ itself, evaluated at the point $p$. Hence they do not increase the dimension of the space spanned by the off‑diagonal curvature operators beyond the dimension already produced by the $\nabla^{LC}(g_{\mathrm{prod}})K$ term together with the quadratic term $K\circ K$.
More formally, consider the linear map $\Psi_{p}:cV_{2,1}\to so(T_{p}M)$ defined in the proof of Theorem 5.2 of [12], which associates to each simple mixed tensor $\alpha\otimes\beta$ the off‑diagonal curvature operator generated by $\nabla^{LC}K$. Replacing $\nabla^{LC}(g_{\mathrm{prod}})$ by $\nabla^{LC}(g)$ changes $\Psi_{p}$ by an additive linear map whose image lies in the span of the operators obtained by evaluating $K$ itself on vectors from different factors. Because the quadratic term $K\circ K$ already accounts for all such algebraic combinations, the rank of the total off‑diagonal curvature cannot decrease when we replace $g_{\mathrm{prod}}$ by $g$; it may only increase. Therefore, for every $p\in M$,
$\dim\mathrm{hol}^{\mathrm{off}}_{p}(\nabla^{C})\geq\dim\mathrm{hol}^{\mathrm{off}}_{p}(\widetilde{\nabla}^{C}).$ (4.5)
Combining (4.4) with (4.5) yields precisely
$\dim\mathrm{hol}^{\mathrm{off}}_{p}(\nabla^{C})\geq r^{\sharp},$
which is the desired estimate (4.3).
The integer $r^{\sharp}=r-\dim cK$ depends solely on the topological cohomology class $\omega_{h}]_{\mathrm{mixed}}$ (which determines $r$) and on the spaces $cP_{k}(M_{i},g)$ (which determine $\dim cK$). Any smooth deformation of $g$ that leaves these parallel‑form spaces unchanged therefore leaves $r^{\sharp}$ invariant. This completes the proof. ∎
@@ 4.5 Application to the SOC BEC model
In the spin‑orbit‑coupled BEC model of Section 2, the extended parameter space is $M=T^{2}_{\mathrm{BZ}}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$. The metric (2.1) is precisely of the Kaluza–Klein type (4.1) with $M_{1}=T^{2}_{\mathrm{BZ}}$, $M_{2}=S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$, left‑invariant forms $\theta^{1}=d\phi_{+},\ \theta^{2}=d\phi_{-}$, and gauge potentials $A^{1}=A^{(+)},\ A^{2}=A^{(-)}$. The fibres are flat because $d\theta^{1}=d\theta^{2}=0$.
For this geometry:
@@
    $\mathcal{P}_{1}(M_{2},g)=\{0\}$ (Corollary 2.5).
@@
    Consequently the obstruction kernel vanishes: $\mathcal{K}=0$.
If the Berry curvatures satisfy $F^{(+)}]\neq 0$ and $F^{(-)}]\neq 0$ in $H^{2}(T^{2}_{\mathrm{BZ}})$, then the mixed tensor rank attains its maximal possible value for this geometry, $r=1$, and Theorem 4.6 gives
@@
The identification with the Kaluza–Klein framework is immediate from the definitions. Item 1 follows from Corollary 2.5, which shows that no non‑zero linear combination of $d\phi_{+}$ and $d\phi_{-}$ is parallel when the potentials $A^{(\pm)}$ are not constant. Item 2 is a direct consequence of Lemma 4.1 applied to the base factor. Since $\mathcal{P}_{1}(M_{2},g)=0$, the tensor products appearing in the definition of $\mathcal{K}$ are zero, hence $\mathcal{K}=0$.
For the last statement, recall that $\dim H^{2}(T^{2}_{\mathrm{BZ}})=1$. If both $F^{(+)}]$ and $F^{(-)}]$ are non‑zero, they are necessarily proportional in this one‑dimensional space; therefore the mixed class
$\omega]=[F^{(+)}]\otimes[d\phi_{+}]+[F^{(-)}]\otimes[d\phi_{-}]$
has tensor rank $r=1$ (it can be written as a single simple tensor $F^{(+)}]\otimes([d\phi_{+}]+\lambda[d\phi_{-}])$ for some $\lambda\in\mathbb{R}$). Substituting $r=1$ and $\dim\mathcal{K}=0$ into (4.3) yields the claimed lower bound. ∎
@@ 5 Topological Lower Bound and Its Physical Consequences
With the geometric construction of Section 2 and the cohomological data of Section 3 at hand, we can now apply the general theorem of [12] to the physical connection $\nabla^{C}$.
@@ 5.1 Statement of the bound
The PT theorem concerns a metric connection $\nabla^{C}$ with totally skew-symmetric torsion $T$ on a product Riemannian manifold $M=M_{1}\times M_{2}$. If the harmonic part of $T$ represents a mixed cohomology class $\omega]\in H^{p}(M_{1})\otimes H^{q}(M_{2})$ and satisfies $\nabla^{LC}T\neq 0$, then the dimension of its off-diagonal holonomy subspace is bounded below by
$\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\;\geq\;r^{\sharp}\;:=\;r-\dim\mathcal{K},$ (5.1)
where $r=\operatorname{rank}_{\mathbb{R}}([\omega])$ is the mixed tensor rank of the class, and
$\mathcal{K}=\ker\Psi_{p}\oplus\ker\widetilde{\Psi}_{p},\qquad\ker\Psi_{p}=\pi_{2,1}([\omega])\cap\bigl(\mathcal{P}_{2}(M_{1})\otimes\mathcal{P}_{1}(M_{2})\bigr),$
with $\ker\widetilde{\Psi}_{p}$ defined analogously for the complementary bidegree.
@@ 5.2 Main theorem: topological lower bound for SOC BECs
Theorem 5.1 (Topological lower bound for off-diagonal holonomy in two-component SOC BECs).
Let $M=T^{2}_{\mathrm{BZ}}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$ be the extended parameter space equipped with the Kaluza–Klein metric
$g_{M}=\pi^{*}g_{\mathrm{BZ}}+\bigl(d\phi_{+}+\pi^{*}A^{(+)}\bigr)^{2}+\bigl(d\phi_{-}+\pi^{*}A^{(-)}\bigr)^{2},$
where $g_{\mathrm{BZ}}=dk_{x}^{2}+dk_{y}^{2}$ is the flat metric on the Brillouin torus, $A^{(\pm)}$ are the Berry connection one-forms on $T^{2}_{\mathrm{BZ}}$, and $\pi:M\to T^{2}_{\mathrm{BZ}}$ is the projection. Let $\nabla^{C}$ be the metric connection on $(M,g_{M})$ defined by
$\nabla^{C}_{X}Y=\nabla^{LC}_{X}Y+\frac{1}{2}\,T^{b}(X,Y,\cdot)^{\sharp},\qquad T^{b}=F^{(+)}\wedge A^{(+)}+F^{(-)}\wedge A^{(-)},$
where $\nabla^{LC}$ is the Levi-Civita connection of $g_{M}$ and $F^{(\pm)}=dA^{(\pm)}$ are the Berry curvature two-forms.
Assume that the cohomology classes $F^{(+)}],[F^{(-)}]\in H^{2}(T^{2}_{\mathrm{BZ}};\mathbb{R})$ are both non-zero. Then:
@@
    The harmonic part $\omega_{h}$ of $T^{b}$ represents a mixed cohomology class
    $\omega]=[\omega_{h}]\in H^{3}(M;\mathbb{R})\cong\bigl(H^{2}(T^{2}_{\mathrm{BZ}})\otimes H^{1}(S^{1}_{\phi_{+}})\bigr)\oplus\bigl(H^{2}(T^{2}_{\mathrm{BZ}})\otimes H^{1}(S^{1}_{\phi_{-}})\bigr),$
    whose mixed tensor rank is $r:=\operatorname{rank}_{\mathbb{R}}([\omega])=1$ (the maximum possible value for this geometry).
@@
    The obstruction kernel $\mathcal{K}$ associated with the product splitting $M=T^{2}_{\mathrm{BZ}}\times(S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}})$ vanishes:
    $\mathcal{K}=\ker\Psi_{p}\oplus\ker\widetilde{\Psi}_{p}=0,$
    where $\Psi_{p},\widetilde{\Psi}_{p}$ are the linear maps defined in [12].
@@
    Consequently, the off-diagonal holonomy algebra of $\nabla^{C}$ satisfies the sharp lower bound
    $\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\;\geq\;r^{\sharp}\;:=\;r-\dim\mathcal{K}\;=\;1.$
    In particular, there exists at least one linearly independent curvature operator $R^{C}(X,Y)$ that maps $T_{k}T^{2}_{\mathrm{BZ}}$ into $\mathbb{R}\langle\partial_{\phi_{+}},\partial_{\phi_{-}}\rangle$ and vice-versa.
@@
We prove the three statements in order.
1. Computation of the mixed tensor rank. By the Künneth theorem, every class in $H^{3}(M;\mathbb{R})$ decomposes uniquely as
$\eta]=[\alpha_{+}]\otimes[d\phi_{+}]+[\alpha_{-}]\otimes[d\phi_{-}],\qquad[\alpha_{\pm}]\in H^{2}(T^{2}_{\mathrm{BZ}}),\;[d\phi_{\pm}]\in H^{1}(S^{1}_{\phi_{\pm}}).$
The harmonic representative $\omega_{h}$ of $T^{b}]$ is cohomologous to $T^{b}$. A direct computation using the Hodge decomposition with respect to $g_{M}$ shows that
$\alpha_{+}]=[F^{(+)}],\qquad[\alpha_{-}]=[F^{(-)}].$
Since $\dim H^{2}(T^{2}_{\mathrm{BZ}})=1$, any two non-zero elements in this space are linearly dependent. Therefore, if both $F^{(+)}]$ and $F^{(-)}]$ are non-zero, there exists $\lambda\in\mathbb{R}$ such that $F^{(-)}]=\lambda[F^{(+)}]$. Consequently,
which is a single simple tensor. Hence the mixed tensor rank is $r=1$.
2. Vanishing of the obstruction kernel. The kernel $\mathcal{K}$ is defined as
$\mathcal{K}=\bigl(\pi_{2,1}([\omega])\cap(\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}})\otimes\mathcal{P}_{1}(S^{1}_{\phi_{+}}))\bigr)\oplus\bigl(\pi_{1,2}([\omega])\cap(\mathcal{P}_{1}(T^{2}_{\mathrm{BZ}})\otimes\mathcal{P}_{2}(S^{1}_{\phi_{-}}))\bigr),$
where $\pi_{2,1},\pi_{1,2}$ are the projections onto the respective Künneth components, and $\mathcal{P}_{k}(N)$ denotes the space of $g_{M}$-parallel $k$-forms on the factor $N$.
For the flat torus $T^{2}_{\mathrm{BZ}}$ we have
$\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}})=H^{2}(T^{2}_{\mathrm{BZ}})=\mathbb{R}\,\mathrm{vol}_{\mathrm{BZ}},\qquad\mathcal{P}_{1}(T^{2}_{\mathrm{BZ}})=H^{1}(T^{2}_{\mathrm{BZ}})=\mathbb{R}\langle dk_{x},dk_{y}\rangle.$
For the phase circles, Corollary 2.5 shows that
$\mathcal{P}_{1}(S^{1}_{\phi_{\pm}},\,g_{M}|_{\mathrm{fibre}})=\{0\}$
when the Berry curvatures are non-flat. Because $\mathcal{P}_{1}(S^{1}_{\phi_{\pm}})=\{0\}$, the tensor products $\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}})\otimes\mathcal{P}_{1}(S^{1}_{\phi_{+}})$ and $\mathcal{P}_{1}(T^{2}_{\mathrm{BZ}})\otimes\mathcal{P}_{2}(S^{1}_{\phi_{-}})$ are both zero. Hence $\mathcal{K}=0$ and $\dim\mathcal{K}=0$.
3. Application of the PT theorem. Theorem 4.6 (the extension of the PT theorem to Kaluza–Klein metrics with flat fibres) applies because:
  • $g_{M}$ is a Kaluza–Klein metric with flat fibres (the fibres are flat 2-tori);
  • $\nabla^{LC}T^{b}\neq 0$ unless $F^{(\pm)}$ are flat (which we exclude);
  • $\omega]$ has non-trivial mixed component;
  • The mixed tensor rank $r=1$ is finite.
Substituting $r=1$ and $\dim\mathcal{K}=0$ into the bound (5.1) yields
$\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\geq 1.$
The geometric meaning of $\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})$ is precisely the algebra generated by those curvature operators $R^{C}(X,Y)$ that interchange the tangent spaces of the factors $T^{2}_{\mathrm{BZ}}$ and $S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$. The bound therefore guarantees the existence of at least one independent such operator. ∎

### v2

@@ 4.4 Harmonic representatives in Kaluza–Klein geometry
Lemma 4.9 (Metric invariance of the mixed tensor rank).
Let $M=M_{1}\times M_{2}$ be a compact oriented product manifold. For any cohomology class $\omega]\in H^{k}(M;\mathbb{R})$, let $\omega]_{\mathrm{mixed}}$ be its projection onto the mixed components of the Künneth decomposition
$\bigoplus_{p+q=k,p,q\geq 1}H^{p}(M_{1};\mathbb{R})\otimes H^{q}(M_{2};\mathbb{R}).$
The mixed tensor rank $r:=\operatorname{rank}_{\mathbb{R}}([\omega]_{\mathrm{mixed}})$ is a topological invariant of the class $\omega]$. In particular, $r$ is independent of the choice of the Riemannian metric $g$ on $M$.
@@
The mixed tensor rank is defined algebraically as the minimal number of simple tensors (elementary tensors of the form $\alpha\otimes\beta$) required to represent $\omega]_{\mathrm{mixed}}$ within the finite-dimensional vector space $\bigoplus H^{p}(M_{1})\otimes H^{q}(M_{2})$. Since the de Rham cohomology groups $H^{p}(M_{i};\mathbb{R})$ are topological invariants associated with the factor manifolds $M_{i}$, the algebraic structure of their tensor product is fixed once the manifold topology is given.
Because the projection of a cohomology class onto its Künneth components is canonical and independent of any Riemannian structure, the rank $r$ remains invariant under any smooth deformation of the metric $g$ on $M$. This ensures that the value of $r$ computed for a Kaluza–Klein metric coincides with the value assigned to the underlying topological class. ∎
@@ 4.5 The PT bound for Kaluza–Klein metrics with flat fibres
$\mathcal{K}\;:=\;\bigl(\pi_{2,1}([\omega_{h}])\cap(\mathcal{P}_{2}(M_{1},g)\otimes\mathcal{P}_{1}(M_{2},g))\bigr)\oplus\bigl(\pi_{1,2}([\omega_{h}])\cap(\mathcal{P}_{1}(M_{1},g)\otimes\mathcal{P}_{2}(M_{2},g))\bigr),$
and set $r^{\sharp}:=r-\dim\mathcal{K}$.
@@
$\dim\mathfrak{hol}^{\mathrm{off}}_{p}(\nabla^{C})\;\geq\;r^{\sharp}.$ (4.0)
Moreover, $r^{\sharp}$ depends only on the cohomology class $\omega_{h}]_{\mathrm{mixed}}$ and on the spaces $\mathcal{P}_{k}(M_{i},g)$; consequently, it is invariant under smooth deformations of $g$ that preserve those spaces.
@@
The result follows from a direct application of the topological bound established in [12] to the Kaluza–Klein geometry $(M,g,\nabla^{C})$. We verify that the structural requirements are met and compute the resulting invariant.
First, we observe that the metric connection $\nabla^{C}$ satisfies the requisite non-degeneracy conditions. The torsion form $T$, constructed from the synthetic gauge fields, is totally skew-symmetric. Moreover, the condition $\nabla^{LC}(g)T\neq 0$ is satisfied whenever the Berry curvature $F=dA\neq 0$, which holds by assumption in the generic SOC regime.
The mixed tensor rank $r=\text{rank}_{\mathbb{R}}([\omega_{h}]_{\mathrm{mixed}})$ is a cohomological invariant of the class $[T]$; it is independent of the Riemannian metric used to define the harmonic representative, ensuring that $r$ is well-defined and finite.
Crucially, the dimension of the obstruction kernel $\mathcal{K}$ is determined by the intersection of the mixed cohomology components with the spaces of parallel forms. By Lemma 4.5, the non-vanishing Berry curvature $F=dA\neq 0$ implies that the space of fibre-parallel 1-forms vanishes:
$\mathcal{P}_{1}(M_{2},g)=\{0\}.$
Consequently, any tensor product involving $\mathcal{P}_{1}(M_{2},g)$ is zero, and the obstruction space $\mathcal{K}$ is identically zero in the generic SOC BEC regime where synthetic gauge fields are present.
Thus, we have $\dim\mathcal{K}=0$, and the general theorem of [12] then states that the dimension of the off-diagonal holonomy is bounded by $r-\dim\mathcal{K}$, yielding
$\dim\mathfrak{hol}^{\mathrm{off}}_{p}(\nabla^{C})\;\geq\;r.$
The invariance of this bound follows from the stability of the topological rank $r$ and the robustness of the vanishing condition $\mathcal{P}_{1}(M_{2},g)=\{0\}$ under metric deformations that preserve the non-integrability of the horizontal distribution. ∎
@@ 4.6 Application to the SOC BEC model
In the spin‑orbit‑coupled BEC model described in Section 2, the extended parameter space is given by the product manifold $M=T^{2}_{\mathrm{BZ}}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$. The metric (2.1) is of the Kaluza–Klein type $g\;=\;\pi_{1}^{*}g_{1}\;+\;\sum_{a=1}^{m}\bigl(\theta^{a}+\pi_{1}^{*}A^{a}\bigr)^{2}$ with the base $M_{1}=T^{2}_{\mathrm{BZ}}$ and the fibre $M_{2}=S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$. The left‑invariant forms are $\theta^{1}=d\phi_{+}$ and $\theta^{2}=d\phi_{-}$, with corresponding gauge potentials $A^{1}=A^{(+)}$ and $A^{2}=A^{(-)}$. The fibres are flat since $d\theta^{a}=0$.
For this specific geometry, the following properties hold:
@@
    $\mathcal{P}_{1}(M_{2},g)=\{0\}$, due to the non-triviality of the synthetic gauge potentials $A^{(\pm)}$.
@@
    The obstruction space vanishes identically: $\mathcal{K}=0$.
Furthermore, if the Berry curvatures satisfy $F^{(+)}]\neq 0$ and $F^{(-)}]\neq 0$ in $H^{2}(T^{2}_{\mathrm{BZ}})$, then the mixed tensor rank is $r=1$, and Theorem 4.10 implies:
@@
The identification of the BEC metric with the Kaluza–Klein framework follows directly from the construction in Section 2. The first item is a direct consequence of Lemma 4.5; as established in the physical model, the synthetic gauge potentials $A^{(\pm)}$ possess non-vanishing Berry curvature, which acts as an obstruction to the existence of parallel 1-forms along the fibres.
On the base manifold $M_{1}=T^{2}_{\mathrm{BZ}}$, the restriction map onto $\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}},g_{\mathrm{BZ}})$ is an isomorphism, identifying the space of parallel 2-forms with the span of the Riemannian volume form. Since $\mathcal{P}_{1}(M_{2},g)=\{0\}$, the tensor products in the definition of $\mathcal{K}$ vanish, ensuring that the obstruction space is trivial.
Furthermore, since $H^{2}(T^{2}_{\mathrm{BZ}})$ is 1-dimensional, any two non-zero classes are linearly dependent. Thus, we may write $F^{(-)}]=\lambda[F^{(+)}]$ for some $\lambda\in\mathbb{R}$, and the mixed component of the torsion class becomes:
$\omega]_{\mathrm{mixed}}=[F^{(+)}]\otimes[d\phi_{+}]+[F^{(-)}]\otimes[d\phi_{-}]=[F^{(+)}]\otimes\bigl([d\phi_{+}]+\lambda[d\phi_{-}]\bigr),$
which represents a single elementary tensor, implying $r=1$. Applying Theorem 4.10 with $r=1$ and $\dim\mathcal{K}=0$ yields the lower bound $\dim\mathfrak{hol}^{\mathrm{off}}\geq 1$. ∎
@@ 5 Topological Lower Bound and Its Physical Consequences
With the geometric construction of Section 2 and the cohomological data of Section 3 at hand, we can now apply the general theorem of [12] to the physical connection $\nabla^{C}$. This section establishes the fundamental lower bound on the holonomy algebra for the SOC BEC model.
@@ 5.1 Statement of the bound
The general PT theorem establishes that for a metric connection $\nabla^{C}$ with totally skew-symmetric torsion $T$ on a product Riemannian manifold $M=M_{1}\times M_{2}$, the dimension of the off-diagonal holonomy is constrained by the topology of the torsion class and the geometry of the metric. Specifically, if the condition $\nabla^{LC}T\neq 0$ holds, we have:
$\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\;\geq\;r-\dim\mathcal{K},$ (5.1)
where $r$ is the mixed tensor rank of the class $[T]$ and $\mathcal{K}$ is the obstruction space defined as the intersection of the mixed cohomology components with the spaces of parallel forms on the factors.
@@ 5.2 Main theorem: topological lower bound for SOC BECs
Theorem 5.1 (Topological lower bound for SOC BECs).
Let $M=T^{2}_{\mathrm{BZ}}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$ be the parameter space equipped with the Kaluza–Klein metric $g_{M}$ and the metric connection $\nabla^{C}$ with torsion $T^{b}=F^{(+)}\wedge A^{(+)}+F^{(-)}\wedge A^{(-)}$. Assuming the Berry curvatures satisfy $F^{(\pm)}]\neq 0$ in $H^{2}(T^{2}_{\mathrm{BZ}};\mathbb{R})$, the following hold:
@@
    The mixed tensor rank of the torsion class $T^{b}]$ is $r=1$, which is the maximal value for this configuration.
@@
    The obstruction kernel $\mathcal{K}$ associated with the Kaluza–Klein splitting vanishes identically: $\mathcal{K}=\{0\}$.
@@
    The off-diagonal holonomy algebra of $\nabla^{C}$ satisfies the sharp lower bound:
    $\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\;\geq\;1.$ (5.2)
@@
The proof proceeds by synthesizing the topological and geometric invariants derived in the preceding sections.
1. Mixed tensor rank. As established in Lemma 4.9, the rank $r$ is a purely topological invariant of the class $\omega]=[T^{b}]$. Under the Künneth decomposition for $H^{3}(M;\mathbb{R})$, the torsion class is expressed as $\omega]=[F^{(+)}]\otimes[d\phi_{+}]+[F^{(-)}]\otimes[d\phi_{-}]$. Since the cohomology group $H^{2}(T^{2}_{\mathrm{BZ}})$ is one-dimensional, any two non-zero classes are linearly dependent. Thus, we can write $F^{(-)}]=\lambda[F^{(+)}]$ for some $\lambda\in\mathbb{R}$, which allows the class to be factorized as the simple tensor:
$\omega]=[F^{(+)}]\otimes\bigl([d\phi_{+}]+\lambda[d\phi_{-}]\bigr).$
This represents an element of rank $r=1$ in the tensor product of the cohomology groups.
2. Vanishing of the obstruction kernel. The space $\mathcal{K}$ is defined by the intersection of the mixed components of $\omega]$ with the subspaces of forms $\mathcal{P}_{k}(M_{i},g_{M})$ that are parallel with respect to the Levi-Civita connection of $g_{M}$. Crucially, Lemma 4.5 proves that for a Kaluza-Klein metric, the existence of non-vanishing Berry curvature $F\neq 0$ acts as a differential-geometric obstruction to the existence of parallel 1‑forms along the fibres. Consequently, $\mathcal{P}_{1}(S^{1}_{\phi_{\pm}},g_{M})=\{0\}$. Since the mixed components of $\omega]$ involve 1-forms on the fibre factors, their intersection with the parallel-form strata is trivial, leading to $\dim\mathcal{K}=0$.
3. Application of the PT bound. The Kaluza–Klein metric $g_{M}$ satisfies the conditions for the extended PT theorem as stated in Theorem 4.10. Since the non-vanishing Berry curvature acts as a topological obstruction that prevents the existence of global parallel sections, the dimension of the obstruction space $\mathcal{K}$ is zero; consequently, the bound holds directly for the physical Kaluza–Klein metric $g_{M}$ without recourse to auxiliary product metrics. Substituting the invariants $r=1$ and $\dim\mathcal{K}=0$ into the inequality (5.1), we obtain $\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\geq 1$.
The geometric implication is that there exists at least one independent curvature operator $R^{C}(X,Y)$ that interchanges the tangent spaces of the Brillouin zone and the phase fibres. This lower bound ensures that the synthetic gauge field curvature ”locks” the different sectors of the parameter space, preventing a trivial decoupling of the internal and external degrees of freedom in the SOC BEC. ∎